# Multi-Field Dilaton Screening Beyond the Thin-Shell Mechanism

### 14 Mar 2026 v1

@@ IV.1 The models
The purpose of this section is to analyse the case of a stringy dilaton, by which we refer to the environmentally dependent (runaway) dilaton scenarios of [51, 27] (see also the screening review [38]). In the Einstein frame, a conformal coupling $\tilde{g}_{\mu\nu}=A^{2}(\chi)\,g_{\mu\nu}$ introduced in section˜III implies that non-relativistic matter acts as a $\chi$-dependent source, so static configurations can be organised by an effective potential composed of both matter and vacuum contributions [38], as in eq.˜17. For the toy constant-coupling choice
@@ V.1 Axions and their effects
Figure 2: Radial profiles (First: dilaton, Second: axion, Third: density, Fourth: dilaton charge) for different axion boundary jumps $\Delta{\mathfrak{a}}={\mathfrak{a}}_{+}-{\mathfrak{a}}_{-}$ and a wide axion gradient using the flux conserving treatment for the axion described in Appendix˜B instead of solving the full equations of motion, with $\ell_{\mathfrak{a}}=0.15$. Here we took $\beta=0.2$, $\zeta=\sqrt{2}$ and $W_{0}=1$. We also took boundary conditions for the dilaton field $\chi^{\prime}(0)=0$, $\chi(0)=\chi_{c}$ corresponding to the pinned regime. Crucially for numerical stability in the presence of exponential factors of the dilaton field, we choose an Earth–like density $\rho_{c}\simeq 10^{-9}M_{\rm pl}^{2}/R_{s}^{2}$, with $V_{0}=0.5\,\rho_{c}$ and $\rho_{\rm env}=0.1\,\rho_{c}$ (an unrealistically small hierarchy).
@@
Figure 3: Radial profiles for the same parameters as in Figure˜2 except for $\rho_{\rm env}=10^{-8}\rho_{c}$ and a star-like object density profile described in Appendix˜B. In this case we solve the full multi-field equations using the microphysical parameters $\Lambda_{\mathfrak{a}}=3\times 10^{-6}M_{\rm pl}$ and $m_{\mathfrak{a}}=3\times 10^{-16}\,\text{eV}$.
@@ VI.1 Pinned gradient suppression
Figure 5: Field profiles and dilaton charge evaluated for a pinned scalar in the presence of a suppressing coupling to the axion in a stellar density profile outlined in Appendix˜B. Here we took $W_{0}=1$, $V_{0}=10^{-5}\rho_{c}$ and $\rho_{\rm env}=10^{-15}\rho_{c}$. We solve the full coupled two-field system using the microphysical axion potentials eq.˜36 and eq.˜35 with $\Lambda_{\mathfrak{a}}=3\times 10^{-7}M_{\rm pl}$ and $m_{\mathfrak{a}}=3\times 10^{-16}\,\text{eV}$.
@@ VI.5 Energy minimisation and suppression of the dilaton charge
The derivation of eq.˜67 assumes that the dominant $\chi_{s}$ dependence of the total energy arises from the axion-driven surface contribution, while the implicit dependence entering through the interior value $\chi_{c}$ is subleading. The validity of this approximation can be tested directly by performing a full numerical energy minimisation of the coupled system, the methodology of which is discussed in Appendix˜B. This is shown in Figure˜9, where the total energy section˜VI.4 is evaluated numerically as a function of $\chi_{s}$. The analytic prediction eq.˜67 reproduces the location of the numerical minimum to within $\lesssim 10^{-11}M_{\rm pl}$ across the range of axion ramp widths considered. We can see that the approximation is best for thinner axion gradients and becomes inaccurate as one considers more gradual axion ramps, going beyond the fast transition window approximation made in the analytic derivation. This is because the dominant energy contribution becomes no-longer localised within a small shell at the surface, as was assumed in eq.˜62, and one needs to include additional corrections from contributions outside this shell width as the axion gradient widens. We have further verified that this residual offset produces no observable change in the inferred charge suppression, and the resulting $L/L_{0}$ (and hence $\beta_{\rm eff}$) remains unchanged at the level relevant for the screening conclusions of this work.
@@
In the BBQ regime the system achieves this by axion backreaction. Because the axion gradient depends on the local value of $W(\chi)$, the coupled equations realise a self–regulating feedback, where shifting $\chi$ reshapes the axion contribution to the dilaton equation. For a sufficiently large hierarchy between the axion mass inside and outside the source (so that the axion adiabatically tracks its effective potential; see Appendix˜B), one finds ${\mathfrak{a}}^{\prime}(r)\propto 1/W^{2}(\chi)$. In the BBQ regime $W(\chi)$ is monotonic with $W_{,\chi}<0$ over the relevant field range, so driving the dilaton to smaller $\chi$ increases $W(\chi)$. In the adiabatic regime the axion gradient is then dynamically suppressed. This prevents the axion contribution from overshooting the matter and potential terms in eq.˜23. If it begins to dominate and pushes $\chi$ downwards, the resulting increase in $W$ reduces $\mathfrak{a}$$^{\prime}$, weakening the axion backreaction and stabilising the system near the value of $\chi_{s}$ for which the net sourcing in the rolling region is minimised. This is why the dilaton profile is ruthlessly flattened and the charge is largely suppressed in Figure˜10 and Figure˜11.
@@
Figure 10: Radial profiles (in units of $r/R_{s}$) illustrating the energy–minimisation (BBQ) regime, for different axion boundary jumps $\Delta{\mathfrak{a}}=0.04$. Here we took $W_{0}=1$, $\beta=0.2$, $\zeta=-2\times 10^{10}$. This is using the numerical energy minimisation scheme to determine $\chi_{s}$, outlined in Appendix˜B.
@@
Figure 12: Radial profiles in the presence of a stellar mass density (see Appendix˜B for details) illustrating the energy–minimisation (BBQ) regime, for an axion boundary jump $\Delta{\mathfrak{a}}=4\times 10^{-9}$. Here we took $W_{0}=1$, $\beta=0.2$, $\zeta=-2\times 10^{10}$. This is using the numerical energy minimisation scheme to determine $\chi_{s}$, also outlined in Appendix˜B.
@@ Appendix A Quadratic kinetic prefactor
The qualitative behaviour of the quadratic–$W$ solutions is shown in Figure˜13 and is most transparently understood in the flux–conserving formulation used in the numerical methods outlined in Appendix˜B. Assuming there is a large hierarchy of the axion masses interior and exterior to the source (as well as the axion being heavy in both environments) so that it tracks the minimum of its effective potential, in spherical symmetry the axion equation implies a conserved radial flux $J$, so that
@@ Appendix B Numerical methodology
For sharply localised surface features (in particular the near–top–hat profiles used in appendix˜B), the coupled system can become numerically stiff in the vicinity of $r\simeq R_{s}$ and can easily enter into non-physical runaway regimes. In these runs the dilaton does not reliably relax to a well–defined exterior minimum within the numerical domain, so there is no stable asymptotic condition that would select a unique boundary solution for $\chi$. We therefore evolve the dilaton equation as an initial–value problem (IVP), imposing regularity at a small inner radius $r_{\rm in}\ll R_{s}$,

### 19 Mar 2026 v2

@@ IV.1 The models
The purpose of this section is to analyse the case of a stringy dilaton, by which we refer to the environmentally dependent (runaway) dilaton scenarios of [51, 27] (see also the screening review [38]). In the Einstein frame, a conformal coupling $\tilde{g}_{\mu\nu}=A^{2}(\chi)\,g_{\mu\nu}$ introduced in Sec.˜III implies that non-relativistic matter acts as a $\chi$-dependent source, so static configurations can be organised by an effective potential composed of both matter and vacuum contributions [38], as in eq.˜17. For the toy constant-coupling choice
@@ V.1 Axions and their effects
Figure 2: Radial profiles (First: dilaton, Second: axion, Third: density, Fourth: dilaton charge) for different axion boundary jumps $\Delta{\mathfrak{a}}={\mathfrak{a}}_{+}-{\mathfrak{a}}_{-}$ and a wide axion gradient using the flux conserving treatment for the axion described in Appendix B instead of solving the full equations of motion, with $\ell_{\mathfrak{a}}=0.15$. Here we took $\beta=0.2$, $\zeta=\sqrt{2}$ and $W_{0}=1$. We also took boundary conditions for the dilaton field $\chi^{\prime}(0)=0$, $\chi(0)=\chi_{c}$ corresponding to the pinned regime. Crucially for numerical stability in the presence of exponential factors of the dilaton field, we choose an Earth–like density $\rho_{c}\simeq 10^{-9}M_{\rm pl}^{2}/R_{s}^{2}$, with $V_{0}=0.5\,\rho_{c}$ and $\rho_{\rm env}=0.1\,\rho_{c}$ (an unrealistically small hierarchy).
@@
Figure 3: Radial profiles for the same parameters as in Figure˜2 except for $\rho_{\rm env}=10^{-8}\rho_{c}$ and a star-like object density profile described in Appendix B. In this case we solve the full multi-field equations using the microphysical parameters $\Lambda_{\mathfrak{a}}=3\times 10^{-6}M_{\rm pl}$ and $m_{\mathfrak{a}}=3\times 10^{-16}\,\text{eV}$.
@@ VI.1 Pinned gradient suppression
Figure 5: Field profiles and dilaton charge evaluated for a pinned scalar in the presence of a suppressing coupling to the axion in a stellar density profile outlined in Appendix B. Here we took $W_{0}=1$, $V_{0}=10^{-5}\rho_{c}$ and $\rho_{\rm env}=10^{-15}\rho_{c}$. We solve the full coupled two-field system using the microphysical axion potentials eq.˜36 and eq.˜35 with $\Lambda_{\mathfrak{a}}=3\times 10^{-7}M_{\rm pl}$ and $m_{\mathfrak{a}}=3\times 10^{-16}\,\text{eV}$.
@@ VI.5 Energy minimisation and suppression of the dilaton charge
The derivation of eq.˜67 assumes that the dominant $\chi_{s}$ dependence of the total energy arises from the axion-driven surface contribution, while the implicit dependence entering through the interior value $\chi_{c}$ is subleading. The validity of this approximation can be tested directly by performing a full numerical energy minimisation of the coupled system, the methodology of which is discussed in Appendix B. This is shown in Figure˜9, where the total energy Sec.˜VI.4 is evaluated numerically as a function of $\chi_{s}$. The analytic prediction eq.˜67 reproduces the location of the numerical minimum to within $\lesssim 10^{-11}M_{\rm pl}$ across the range of axion ramp widths considered. We can see that the approximation is best for thinner axion gradients and becomes inaccurate as one considers more gradual axion ramps, going beyond the fast transition window approximation made in the analytic derivation. This is because the dominant energy contribution becomes no-longer localised within a small shell at the surface, as was assumed in eq.˜62, and one needs to include additional corrections from contributions outside this shell width as the axion gradient widens. We have further verified that this residual offset produces no observable change in the inferred charge suppression, and the resulting $L/L_{0}$ (and hence $\beta_{\rm eff}$) remains unchanged at the level relevant for the screening conclusions of this work.
@@
In the BBQ regime the system achieves this by axion backreaction. Because the axion gradient depends on the local value of $W(\chi)$, the coupled equations realise a self–regulating feedback, where shifting $\chi$ reshapes the axion contribution to the dilaton equation. For a sufficiently large hierarchy between the axion mass inside and outside the source (so that the axion adiabatically tracks its effective potential; see Appendix B), one finds ${\mathfrak{a}}^{\prime}(r)\propto 1/W^{2}(\chi)$. In the BBQ regime $W(\chi)$ is monotonic with $W_{,\chi}<0$ over the relevant field range, so driving the dilaton to smaller $\chi$ increases $W(\chi)$. In the adiabatic regime the axion gradient is then dynamically suppressed. This prevents the axion contribution from overshooting the matter and potential terms in eq.˜23. If it begins to dominate and pushes $\chi$ downwards, the resulting increase in $W$ reduces $\mathfrak{a}$$^{\prime}$, weakening the axion backreaction and stabilising the system near the value of $\chi_{s}$ for which the net sourcing in the rolling region is minimised. This is why the dilaton profile is ruthlessly flattened and the charge is largely suppressed in Figure˜10 and Figure˜11.
@@
Figure 10: Radial profiles (in units of $r/R_{s}$) illustrating the energy–minimisation (BBQ) regime, for different axion boundary jumps $\Delta{\mathfrak{a}}=0.04$. Here we took $W_{0}=1$, $\beta=0.2$, $\zeta=-2\times 10^{10}$. This is using the numerical energy minimisation scheme to determine $\chi_{s}$, outlined in Appendix B.
@@
Figure 12: Radial profiles in the presence of a stellar mass density (see Appendix B for details) illustrating the energy–minimisation (BBQ) regime, for an axion boundary jump $\Delta{\mathfrak{a}}=4\times 10^{-9}$. Here we took $W_{0}=1$, $\beta=0.2$, $\zeta=-2\times 10^{10}$. This is using the numerical energy minimisation scheme to determine $\chi_{s}$, also outlined in Appendix B.
@@ Appendix A Quadratic kinetic prefactor
The qualitative behaviour of the quadratic–$W$ solutions is shown in Figure˜13 and is most transparently understood in the flux–conserving formulation used in the numerical methods outlined in Appendix B. Assuming there is a large hierarchy of the axion masses interior and exterior to the source (as well as the axion being heavy in both environments) so that it tracks the minimum of its effective potential, in spherical symmetry the axion equation implies a conserved radial flux $J$, so that
@@ Appendix B Numerical methodology
For sharply localised surface features (in particular the near–top–hat profiles used in App.˜B), the coupled system can become numerically stiff in the vicinity of $r\simeq R_{s}$ and can easily enter into non-physical runaway regimes. In these runs the dilaton does not reliably relax to a well–defined exterior minimum within the numerical domain, so there is no stable asymptotic condition that would select a unique boundary solution for $\chi$. We therefore evolve the dilaton equation as an initial–value problem (IVP), imposing regularity at a small inner radius $r_{\rm in}\ll R_{s}$,